Two disks with the same inertia 1kg are rotating along the same axis.
• Disk A has radius 2m and disk B has radius 4m.
• Disk A is rotating with 2rad/s, ω direction up
• Disk B is rotating with 1rad/s, ω direction down
• Now the two disks are jointed together, calculate
the final angular momentum of the combined disks
1
Expert's answer
2020-06-08T10:27:58-0400

As per the given question,

Two disk have the same inertia it means "m_1=m_2=1 kg"

The radius of the disk A "(r_1)=2m"

and radius of the disk B"(r_2) =4m"

angular velocity of the disk A is "(\omega_1)=2rad/sec (upward)"

Angular speed of the disk B is "(\omega_2)=1 rad/sec(downward)"

Two disks are getting joined, let the final angular speed if "\omega"

Applying the conservation of angular momentum,

"I_1\omega_1-I_2\omega_2=(I_1+I_2)\omega"


"\Rightarrow \omega=\frac{I_1\omega_1-I_2\omega_2}{I_1+I_2}"


"\Rightarrow \omega=\frac{m_1r_1^2 \omega_1-m_2r_2^2\omega_2}{m_1 r_1^2+m_2 r_2^2}"


"\Rightarrow \omega=\frac{8-16}{4+16} rad/sec"


"\Rightarrow \omega=\frac{-8}{20}=-0.4 rad/sec (downward)"